OpenStudy (anonymous):

Angela pays $250 in advance on her account at an an athletic club. After four visits, Angela has $222 in her account. After 10 visits, Angela has $180 in her account. If Angela is charged the same amount for each visit, how much money does she have in her account after 17 visits?

OpenStudy (anonymous):

250-17*7 = 250 - 119 = 131

OpenStudy (anonymous):

thank you